DETERMINATION OF THE IONIZATION AND DISSOCIATION ENERGIES OF MOLECULAR HYDRO-GEN, H_{2}

\underline{JINJUN LIU}, URS HOLLENSTEIN, and FRÉDÉRIC MERKT, Laboratorium für Physikalische Chemie, ETH-Zürich, 8093 Zürich, Switzerland; EDCEL J. SALUMBIDES, JEROEN C. J. KOELE-MEIJ, KJELD S. E. EIKEMA, and WIM UBACHS, Laser Centre, Department of Physics and Astronomy, Vrije Universiteit, De Boelelaan 1081, 1081 HV Amsterdam, The Netherlands.

The transition wave number from the EF ^{1}Σ_{g}^{+}(v = 0, N = 1) energy level of ortho-H_{2} to the 54p1_{1}(0) Rydberg state below the X^{+} ^{2}Σ_{g}^{+}(v^{+} = 0, N^{+} = 1) ground state of ortho-H_{2}^{+} has been measured to be 25209.99756±(0.00022)_{statistical}±(0.00007)_{systematic} cm^{−1}. Combining this result with previous experimental and theoretical results for other energy level intervals, the ionization and dissociation energies of the hydrogen molecule have been determined to be 124417.49113(37) cm^{−1} and 36118.06962(37) cm^{−1}, respectively, which represents a precision improvement over previous experimental and theoretical results by more than one order of magnitude. The new value of the ionization energy can be regarded as the most precise and accurate experimental result of this quantity, whereas the dissociation energy is a hybrid experimental-theoretical determination.